# ICAT: Evaluating Completeness of Factual Information in Long-form Text Generation

### v1

@@ Abstract
This paper presents ICAT, an evaluation framework for measuring coverage of diverse factual information in long-form text generation. ICAT breaks down a long output text into a list of atomic claims and not only verifies each claim through retrieval from a (reliable) knowledge source, but also computes the alignment between the atomic factual claims and various aspects expected to be presented in the output. We study three implementations of the ICAT framework, each with a different assumption on the availability of aspects and alignment method. By adopting data from the diversification task in the TREC Web Track and the ClueWeb corpus, we evaluate the ICAT framework. We demonstrate strong correlation with human judgments and provide comprehensive evaluation across multiple state-of-the-art LLMs. Our framework further offers interpretable and fine-grained analysis of diversity and coverage. Its modular design allows for easy adaptation to different domains and datasets, making it a valuable tool for evaluating the qualitative aspects of long-form responses produced by LLMs.
@@
Chris Samarinas1thanks: Equal contribution.    Alexander Krubner211footnotemark: 1thanks: Work done while visiting the Center for Intelligent Information Retrieval at UMass Amherst.    Alireza Salemi1 Youngwoo Kim1    Hamed Zamani1 1University of Massachusetts Amherst  2Salzburg University of Applied Sciences 1{csamarinas, asalemi, youngwookim, zamani}@cs.umass.edu 2akrubner.bin-m2022@fh-salzburg.ac.at
@@ 1 Introduction
Recent advancements in large language model (LLM) development have enabled us to go beyond phrase-level and sentence-level generation and build systems capable of generating fluent and statistically plausible long responses. The evaluation of these systems remains a challenging area of research Samarinas et al. (2024). Numerous works have proposed methods that measure lexical overlap with one or more reference outputs, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005), as well as semantic similarity-based methods, such as BERTScore Zhang et al. (2019) and G-Eval Liu et al. (2023), for long-form text generation. One major issue with all these methods is that many different long outputs by LLMs can be considered equally acceptable; given the size of the output space in long-form text generation, it is impractical to create reference texts that can cover all possible long outputs at scale. This has recently motivated researchers to develop reference-free evaluation methodologies, such as Prism Agrawal et al. (2021); Thompson and Post (2020). However, they are often unable to evaluate hallucinations and biases in LLMs. FActScore Min et al. (2023) and VERISCORE Song et al. (2024) address this issue by looking at fine-grained atomic claims in the generated output that can be backed up by retrieved information from a given (reliable) source. These methods solely evaluate the factual accuracy in the generated output.
We argue that factual accuracy is not sufficient when evaluating LLMs. Imagine a user submits the query “is coffee good for health?” to an LLM. Given the current state of knowledge in medicine and related disciplines, there are some health benefits and also some health risks to consuming coffee. If the LLM generates an output that only covers the health benefits of coffee by making factual and grounded arguments, it would achieve a perfect FActScore and VERISCORE. Therefore, answer completeness and coverage of diverse factual information should be a major component when evaluating LLMs for long-form text generation, such as non-factoid and complex comparative question answering, report generation, etc.
This paper introduces ICAT, [^1] ICAT: Information Coverage & Accuracy in Text. Code is publicly available at https://github.com/algoprog/ICAT a reference-free evaluation framework for long-form text generation that considers both factual accuracy and coverage of diverse factual information. Inspired by FActScore Min et al. (2023) and VERISCORE Song et al. (2024), ICAT breaks down the generated long text into atomic claims. Through retrieval from a (reliable) corpus $C$ or the Web, ICAT verifies each atomic claim to ensure its factuality. To measure completeness and coverage of diverse facts, ICAT requires a set of diverse aspects to compute an alignment between each atomic factual claim in the LLM response and the set of diverse aspects. We study three implementations of the ICAT framework as follows: ICAT-M assumes that a ground-truth set of diverse claims are obtained manually and is available to the evaluation framework. It also assumes that the groundtruth relevance annotation for each document in the corpus $C$ to each aspect is provided. Using this information, the retrieval model can identify which aspect is being covered by each atomic factual claim in the LLM response. ICAT-S similarly assumes that a groundtruth set of diverse claims are obtained manually, however no aspect-level relevance judgment is available. Therefore, it uses an LLM to conduct pseudo-labeling and performing alignment between the atomic factual claims and the set of aspects. On the other hand, ICAT-A assumes that the aspect set is not available, so it first uses an LLM to automatically generate diverse aspects of the input and then conduct pseudo-labeling for alignment, as is done in the second variant.
@@
In our experiments, we use ClueWeb The Lemur Project (2009) as the retrieval corpus. We solely focus on the English documents of the ClueWeb collection. For web-based grounding, we use the Brave Search API. For experiments, we rely on the input topics from the TREC Web Track Clarke et al. (2009, 2010, 2011, 2012). The argument for this decision is based on the fact that TREC Web Track topics have also been used for search result diversification. This means that the topics include up to 7 subtopics (aspects) and documents are provided with aspect-level relevance annotations. Our experiments show that there is strong corelation of ICAT with human judgments (ICAT-S and ICAT-A achieving Pearson’s $\rho>0.4$), showcasing the utility of this framework for evaluating coverage of diverse factual information in LLM responses without human input.
@@
By offering a modular and adaptable framework, ICAT enables researchers to tailor the evaluation process to specific needs, making it a valuable tool for assessing the qualitative aspects of long-form responses produced by LLMs. The decomposition of LLM outputs into atomic claims and their alignment with specific topics makes the evaluation process highly interpretable - evaluators can trace exactly which claims support which topics and identify gaps in coverage. This granular analysis capability, combined with the framework’s ability to evaluate both factual accuracy and topic coverage, provides a more comprehensive assessment compared to existing metrics that only measure one of these aspects. ICAT’s implementation is available at https://github.com/algoprog/ICAT.
@@ 2 Related Work
LLM Evaluation Frameworks Several frameworks have been proposed for evaluating different aspects of LLM performance, including factuality Min et al. (2023); Song et al. (2024) and dialogue quality Mehri and Eskenazi (2020). Concurrent to this research, the AutoNuggetizer framework Pradeep et al. (2024) used LLMs to generate and assess the coverage of nuggets in text. However, in that framework, the nuggets were specific claims rather than higher-level subtopics. Moreover, AutoNuggetizer solely focuses on nugget coverage and does not consider factuality of responses.
@@ 3.3 Aspect Coverage Assessment
3.4 Variants of ICAT
We study three variants of ICAT implementations based on how they obtain query aspects and compute the alignment between atomic claims and aspects. Table 1 describes the approaches used in these three variants and highlights their differences.
@@ 3.4 Variants of ICAT
ICAT-M N/A 0.196 0.184 0.146 0.216 0.204 0.166
ICAT-S Llama-3.1-8B 0.417 0.400 0.314 0.534 0.521 0.442
Llama-3.1-70B 0.489 0.452 0.376 0.612 0.589 0.486
ICAT-A Llama-3.1-8B 0.417 0.400 0.314 0.548 0.529 0.447
Llama-3.1-70B 0.500 0.459 0.382 0.623 0.596 0.492
Table 2: Correlation of information coverage in the proposed evaluation methods and manual human annotations.
@@ 4.2 Topic Generation
Previous works have shown that LLMs can be very effective in query subtopic generation Samarinas et al. (2022). In our framework, for generating ground-truth topics given a query, we use the same base LLM as the one used in claim generation. Here we found that even without fine-tuning, Llama 3.1 8B can produce relevant topics. In order to reduce the need for extra resources to use a base and fine-tuned version of the LLM for claim generation, we use the VLLM library Kwon et al. (2023) to load the base model only once in memory and efficiently serve the adapter for the fine-tuned version.
@@ 4.3 Claim Grounding
We limit the snippet length because both NLI and dense embeddings models based on small pre-trained transformer LMs like BERT tend to have lower performance as the input length increases.
@@ 4.4 Aspect-Claim Alignment
Topic coverage is assessed using the same base LLM with claim and topic generation. Given a query, a list of enumerated atomic claims and a list of ground truth topics, the LLM is prompted to produce a list of covered topic ids with their associated claim ids in structured jsonl format.
ICAT-M ICAT-S ICAT-A
LLM $S_{\text{fact}}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-M}_{1}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-S}_{1}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-A}_{1}$
Corpus-based Retrieval
GPT-4 0.343 0.416 0.327 0.453 0.346 0.463 0.354
Llama-3-70B 0.327 0.451 0.335 0.464 0.355 0.466 0.354
Mixtral-8x22B 0.344 0.370 0.297 0.414 0.342 0.409 0.339
Openchat 3.5 (7B) 0.340 0.413 0.329 0.429 0.348 0.424 0.347
Web-based Retrieval
GPT-4 0.363 0.436 0.347 0.473 0.366 0.483 0.374
Llama-3-70B 0.347 0.471 0.355 0.484 0.375 0.486 0.374
Mixtral-8x22B 0.364 0.390 0.317 0.434 0.362 0.429 0.359
Openchat 3.5 (7B) 0.360 0.433 0.349 0.449 0.368 0.444 0.367
@@ 5 Experimental Setup
We conducted our experiments using the ClueWeb09 Category B corpus–a large-scale web collection with over 50 million English documents The Lemur Project (2009). This corpus has been used in TREC Web Track from 2009 to 2012 , consisting of 200 topics, derived from a commercial search engine’s query log, balanced for popularity. Each topic includes a topic title (i.e., often used as the keyword search query), a description (i.e., detailed description of the information need), type, and subtopics (i.e., diverse aspects of the topic). The relevance judgments encompass 38,637 query-document pairs, with 19.06% (7366) marked as relevant. The dataset’s unique advantage lies in its comprehensive coverage of internet content and human-annotated relevance judgments for topical diversity assessment. Relevance was judged either binary or on a five-point scale (later converted to binary), with documents considered relevant when containing useful information for specific subtopics. In our experiments with this collection, we filtered out spam documents using the Waterloo spam scorer Cormack et al. (2011) with the threshold of 70%. We used BM25 to retrieve 1000 documents for each topic (given its title as the query string) and considered these documents for retrieval in our factual verification process. The query set comprises 50 carefully selected faceted queries, each containing 3-8 subtopics representing different aspects of the information need.
@@ 6 Experimental Results
To validate ICAT’s effectiveness, we conducted a comprehensive human evaluation study using Amazon Mechanical Turk (AMT). For each query-answer pair, three independent annotators assessed the coverage of aspects through a custom interface (Figure 7 in Appendix A.4). We limited the HITs to adult workers from the US, UK, Australia and Ireland, with over 98% approval rate who have completed at least 5,000 assignments. The annotators were tasked with identifying whether specific aspects are present in a given LLM-generated text and highlighting corresponding text evidence for each identified aspect. To ensure quality annotations, we provided detailed guidelines with two reference examples. We use majority voting across annotators. The study achieved substantial inter-annotator agreement with Fleiss’s $\kappa$ = 0.829, which is considered as a substantial agreement.
For each query, we calculated Coverage Scores based on the set of covered topics identified by each evaluation method (ICAT variants) and by human annotators, relative to the set of ground truth topics. These per-query coverage scores were then used to compute linear and rank-bsaed correlation metrics (i.e., Pearson’s $\rho$, Spearman’s $\rho$, and Kendall’s $\tau$) between the automated ICAT methods and human judgments.
The correlation analysis between ICAT variants and human judgments (see Table 2) reveals strong performance across most evaluation methods. Using Llama-3.1-70B as the coverage model, ICAT-A achieves the strongest correlations (Pearson’s $\rho$ = 0.500, p < 0.01; Spearman’s $\rho$ = 0.459, p < 0.01). ICAT-S demonstrates similar performance with the same coverage model (Pearson’s $\rho$ = 0.489, p < 0.01; Spearman’s $\rho$ = 0.452, p < 0.01). While computationally efficient, ICAT-M shows weaker correlation (Pearson’s $\rho$ = 0.196, p < 0.05) with human assessments, highlighting the importance of sophisticated coverage models in improving evaluation accuracy.
@@
Our experimental results reveal distinct patterns in how different LLMs balance factuality and coverage (see Table 3). Llama-3-70B demonstrates superior Coverage Score (0.451 for corpus-based and 0.471 for web-based retrieval), while GPT-4 and Mixtral-8x22B show comparable factuality score (0.343 and 0.344, respectively). However, Mixtral exhibits notably lower Coverage Score 0.370 compared to GPT-4’s 0.416, resulting in lower overall $\text{ICAT}_{1}$ scores (0.297 vs 0.327). Notably, Openchat 3.5, despite being the smallest model evaluated, maintains competitive performance across all metrics ($S_{\text{fact}}$ = 0.340, $S_{\text{coverage}}$ = 0.413, $\text{ICAT}_{1}$ = 0.329). When employing ICAT-S and ICAT-A with Llama-3.1-70B as the coverage model, we observe improved Coverage Scores across all models. Llama-3-70B maintains its leading position for corpus-based retrieval, with the highest overall $\text{ICAT}_{1}$ score of 0.355, followed closely by GPT-4 (i.e., 0.346). The results suggest that using Llama-3.1-70B as the coverage model produces more conservative but potentially more reliable assessments compared to Llama-3.1-8B, as evidenced by more consistent Coverage Scores across evaluated LLMs.

### v4

# Beyond Factual Accuracy: Evaluating Coverage of Diverse Factual Information in Long-form Text Generation

@@ Abstract
This paper presents ICAT,111source code: https://github.com/algoprog/ICAT an evaluation framework for measuring coverage of diverse factual information in long-form text generation. ICAT breaks down a long output text into a list of atomic claims and not only verifies each claim through retrieval from a (reliable) knowledge source, but also computes the alignment between the atomic factual claims and various aspects expected to be presented in the output. We study three implementations of the ICAT framework, each with a different assumption on the availability of aspects and alignment method. By adopting data from the diversification task in the TREC Web Track and the ClueWeb corpus, we evaluate the ICAT framework. We demonstrate strong correlation with human judgments and provide comprehensive evaluation across multiple state-of-the-art LLMs. Our framework further offers interpretable and fine-grained analysis of diversity and coverage. Its modular design allows for easy adaptation to different domains and datasets, making it a valuable tool for evaluating the qualitative aspects of long-form responses produced by LLMs.
@@
Chris Samarinas1thanks: Equal contribution.    Alexander Krubner211footnotemark: 1thanks: Work done while visiting CIIR at UMass Amherst.    Alireza Salemi1 Youngwoo Kim1    Hamed Zamani1 1University of Massachusetts Amherst  2Salzburg University of Applied Sciences 1{csamarinas, asalemi, youngwookim, zamani}@cs.umass.edu 2akrubner.bin-m2022@fh-salzburg.ac.at
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has revolutionized long-form text generation, enabling increasingly sophisticated applications from report writing to complex question answering. However, this progress has highlighted a critical challenge: how do we effectively evaluate not just the factual accuracy of generated content, but also its completeness and coverage of diverse perspectives? While recent work has made significant progress in developing comprehensive evaluation frameworks for LLMs Li et al. (2024), existing metrics often focus on isolated aspects like factuality or response quality Min et al. (2023), failing to capture the multi-dimensional nature of high-quality long-form text.
The evaluation of long-form text generation presents unique challenges that go beyond traditional metrics Samarinas et al. (2024). Lexical overlap metrics like BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005) are fundamentally limited by their reliance on surface-level text similarity, making them inadequate for evaluating semantically equivalent but lexically different expressions. While more recent approaches like BERTScore Zhang et al. (2019) and G-Eval Liu et al. (2023) attempt to address this through semantic similarity, they still face fundamental limitations when applied to long-form content. These metrics cannot effectively verify factual accuracy or assess whether the response comprehensively covers all relevant aspects of a topic. The vast space of possible acceptable outputs makes it impractical to create comprehensive reference texts, leading to the development of reference-free evaluation methodologies like Prism Agrawal et al. (2021); Thompson and Post (2020). However, these approaches often struggle to effectively identify hallucinations and biases.
Recent approaches like FActScore Min et al. (2023) and VERISCORE Song et al. (2024) have addressed the factuality challenge by evaluating atomic claims against reliable sources, but factual accuracy alone is insufficient. A unified metric that considers both factuality and coverage is crucial not only for evaluation but also for optimizing LLM performance. Such a metric could serve as a reward function in reinforcement learning to simultaneously improve both the factual accuracy and comprehensive coverage of LLM outputs.
Consider a user asking “What are the health effects of coffee consumption?” While an LLM response might present entirely factual claims about coffee’s benefits, such as its role in improving alertness and potential protective effects against certain diseases, failing to address known health risks (like anxiety or sleep disruption) would present an incomplete and potentially misleading picture. This illustrates a critical gap in current evaluation approaches: the need to assess not just the accuracy of individual claims, but also the comprehensive coverage of diverse relevant information. This is particularly crucial for applications like medical information systems, policy analysis, or educational content where balanced, complete information is essential for informed decision making.
This paper introduces ICAT, a novel evaluation framework that addresses this gap by measuring both factual accuracy and coverage of diverse factual information in long-form text generation. Our key contributions include: 1) a modular evaluation framework that decomposes long-form text into atomic claims and evaluates both their factual accuracy and their alignment with expected aspects, 2) three implementations with varying degrees of automation, suitable for different evaluation scenarios, and 3) a comprehensive evaluation of various LLMs and demonstrating strong correlation with human judgments in a user study.
ICAT first breaks down the generated long text into atomic claims. Through retrieval from a (reliable) corpus $C$ or the Web, ICAT verifies each atomic claim to ensure its factuality. To measure completeness and coverage of diverse facts, ICAT requires a set of diverse aspects to compute an alignment between each atomic factual claim in the LLM response and the set of diverse aspects. We study three implementations of the ICAT framework with different assumptions on the availability of aspects and alignment methods.
@@
In our experiments, we use ClueWeb The Lemur Project (2009) as the retrieval corpus. We solely focus on the English documents of the ClueWeb collection. In addition to custom open-source retrieval from ClueWeb, we explore web-based grounding using the Brave Search API. For experiments, we rely on the input queries from the TREC Web Track Clarke et al. (2009, 2010, 2011, 2012). The argument for this decision is based on the fact that TREC Web Track queries have also been used for search result diversification. This means that the queries include up to 7 aspects and documents are provided with aspect-level relevance annotations. Our experiments show that there is relatively strong corelation of ICAT with human judgments, showcasing the utility of this framework for evaluating coverage of diverse factual information in LLM responses without human input.
@@
By offering a modular and adaptable framework, ICAT enables researchers to tailor the evaluation process to specific needs, making it a valuable tool for assessing the qualitative aspects of long-form responses produced by LLMs. The decomposition of LLM outputs into atomic claims and their alignment with specific topics makes the evaluation process highly interpretable - evaluators can trace exactly which claims support which topics and identify gaps in coverage. This granular analysis capability, combined with the framework’s ability to evaluate both factual accuracy and topic coverage, provides a more comprehensive assessment compared to existing metrics that only measure one of these aspects.
@@ 2 Related Work
LLM Evaluation Frameworks Several frameworks have been proposed for evaluating different aspects of LLM performance, including factuality Min et al. (2023); Song et al. (2024) and dialogue quality Mehri and Eskenazi (2020). Concurrent to this research, the AutoNuggetizer framework Pradeep et al. (2024) used LLMs to generate and assess the coverage of nuggets in text. The EXAM++ framework Farzi and Dietz (2024) evaluates information coverage by checking if system responses can answer query-related exam questions, focusing on high-level answerability. In contrast, our work provides more fine-grained explainability by aligning individual atomic claims with retrieved evidence to verify both factual accuracy and aspect coverage at a more granular level. This allows us to not only assess if an aspect is covered, but also identify the specific claims and evidence supporting that coverage.
@@ 3 ICAT
Variants of ICAT
We study three variants of ICAT implementations based on how they obtain query aspects and compute the alignment between atomic claims and aspects. Table 1 describes the approaches used in these three variants and highlights their differences. ICAT-M assumes that a ground-truth set of diverse aspects are obtained manually and is available to the evaluation framework. It also assumes that ground-truth relevance annotations for each document in the corpus $C$ to each aspect are provided. Using this information, the retrieval model can identify which aspect is being covered by each atomic factual claim in the LLM response. ICAT-S similarly assumes that a ground-truth set of diverse aspects are obtained manually, however no aspect-level relevance judgment is available. Therefore, it uses an LLM to conduct pseudo-labeling and perform alignment between the atomic factual claims and the set of aspects. ICAT-A assumes that the aspect set is not available, so it first uses an LLM to automatically generate diverse aspects of the input and then conducts pseudo-labeling for alignment.
@@ 3.3 Aspect Coverage Assessment
ICAT-M N/A 0.196 0.184 0.146 - - -
ICAT-S Llama-3.1-8B 0.247 0.255 0.200 0.294 0.252 0.204
Llama-3.1-70B 0.422 0.446 0.368 0.525 0.443 0.358
ICAT-A Llama-3.1-8B 0.161 0.125 0.097 0.033 0.021 0.014
Llama-3.1-70B 0.246 0.218 0.173 0.055 0.054 0.045
Table 2: Correlation of information coverage in the proposed evaluation methods with manual human annotations. Though ICAT-A shows weaker correlation than ICAT-S, we observed that auto-generated topics were more comprehensive and higher quality than the existing ones in the TREC dataset, suggesting ICAT-A would perform better with more complete ground-truth topics.
@@ 4.1 Atomic Claim Generation
Our human evaluation comparing the fine-tuned 8B model with zero-shot 70B model revealed an interesting trade-off: while the 70B model achieves better semantic similarity with ground-truth facts, the fine-tuned 8B model produces better-formulated claims with higher fact precision (0.838 vs 0.753). This superior decontextualization is crucial for accurate retrieval-based grounding, as claims must be self-contained to match against corpus snippets effectively. This finding, combined with the significant computational cost savings (approximately 8.75× less memory and proportionally faster inference), justifies our design choice of using a fine-tuned smaller model. We plan to release a similarly fine-tuned adapter for fact-topic alignment to close the performance gap with the 70B model in that component as well. A detailed evaluation of the atomic claim generation quality is provided in Appendix A.5.
@@ 4.2 Topic Generation
Previous works have shown that LLMs can be very effective in query subtopic generation Samarinas et al. (2022). In our framework, for generating ground-truth topics given a query, we use the same base LLM as the one used in claim generation. Here we found that even without fine-tuning, Llama 3.1 8B can produce relevant topics. In order to reduce the need for extra resources to use a base and fine-tuned version of the LLM for claim generation, we use the VLLM library Kwon et al. (2023) to load the base model only once in memory and efficiently serve the adapter for the fine-tuned version. The quality evaluation of topic generation is detailed in Appendix A.5.
@@ 4.3 Claim Grounding
We limit the snippet length because both NLI and dense embeddings models based on small pre-trained transformer LMs like BERT tend to have lower performance as the input length increases. The choice of using a specialized NLI model rather than a zero-shot LLM for claim verification was driven by both efficiency and effectiveness considerations. Our DeBERTa-based model achieves strong performance across multiple NLI benchmarks (MNLI: 0.903, FEVER-NLI: 0.777, ANLI-all: 0.579) while being orders of magnitude faster and more resource-efficient than LLM-based approaches. Given that grounding requires processing numerous claim-snippet pairs, efficiency is crucial for practical deployment and potential use as a reward function in reinforcement learning optimization of LLMs. A comprehensive evaluation of the NLI model performance across benchmarks is provided in Appendix A.5.
@@ 4.4 Aspect-Claim Alignment
Topic coverage is assessed using the same base LLM with claim and topic generation. Given a query, a list of enumerated atomic claims and a list of ground truth topics, the LLM is prompted to produce a list of covered topic ids with their associated claim ids in structured jsonl format. The accuracy of the aspect-claim alignment module is evaluated in Appendix A.5.
ICAT-M ICAT-S ICAT-A LLM $S_{\text{fact}}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-M}_{1}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-S}_{1}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-A}_{1}$ Corpus-based Retrieval GPT-4 0.343 0.416 0.327 0.453 0.346 0.463 0.354 Llama-3-70B 0.327 0.451 0.335 0.464 0.355 0.466 0.354 Mixtral-8x22B 0.344 0.370 0.297 0.414 0.342 0.409 0.339 Openchat 3.5 (7B) 0.340 0.413 0.329 0.429 0.348 0.424 0.347 Web-based Retrieval GPT-4 0.748 - - 0.551 0.634 0.541 0.627 Llama-3-70B 0.714 - - 0.556 0.625 0.486 0.578 Mixtral-8x22B 0.749 - - 0.503 0.601 0.429 0.545 Openchat 3.5 (7B) 0.741 - - 0.520 0.611 0.444 0.555
@@ 5 Experimental Setup
We conducted our experiments using the ClueWeb09 Category B corpus–a large-scale web collection with over 50 million English documents The Lemur Project (2009). This corpus has been used in TREC Web Track from 2009 to 2012 Clarke et al. (2009, 2010, 2011, 2012), consisting of 200 topics, derived from a commercial search engine’s query log, balanced for popularity. Each topic includes a topic title (i.e., often used as the keyword search query), a description (i.e., detailed description of the information need), type, and subtopics (i.e., diverse aspects of the topic). The relevance judgments encompass 38,637 query-document pairs, with 19.06% (7366) marked as relevant. The dataset’s unique advantage lies in its comprehensive coverage of internet content and human-annotated relevance judgments for topical diversity assessment. Relevance was judged either binary or on a five-point scale (later converted to binary), with documents considered relevant when containing useful information for specific subtopics. In our experiments with this collection, we filtered out spam documents using the Waterloo spam scorer Cormack et al. (2011) with the threshold of 70%. We used BM25 to retrieve 1000 documents for each topic (given its title as the query string) and considered these documents for retrieval in our factual verification process. From the original 200 TREC topics, we filtered out approximately 21 queries that were not suitable for long-form text generation evaluation (e.g., queries asking for images, URLs, or specific data points), resulting in 179 valid queries. For human evaluation and detailed component analysis, we randomly sampled 50 queries from this filtered set due to cost constraints. These 50 queries were also used for LLM evaluation to ensure consistency and enable direct comparison with human judgments.
@@ 6 Experimental Results
To validate ICAT’s effectiveness, we conducted a comprehensive human evaluation study using Amazon Mechanical Turk (AMT). For each query-answer pair, three independent annotators assessed the coverage of aspects through a custom interface (Figure 4 in Appendix A.4). We limited the HITs to adult workers from the US, UK, Australia and Ireland, with over 98% approval rate who have completed at least 5,000 assignments. The annotators were tasked with identifying whether specific aspects are present in a given LLM-generated text and highlighting corresponding text evidence for each identified aspect. To ensure quality annotations, we provided detailed guidelines with two reference examples. We use majority voting across annotators. The study achieved substantial inter-annotator agreement with Fleiss’s $\kappa$ = 0.829, which is considered as a substantial agreement. For each query, we calculated Coverage Scores based on the set of covered topics identified by each evaluation method (ICAT variants) and by human annotators, relative to the set of ground truth topics. These per-query coverage scores were then used to compute linear and rank-bsaed correlation metrics (i.e., Pearson’s $\rho$, Spearman’s $\rho$, and Kendall’s $\tau$) between the automated ICAT methods and human judgments.
The correlation analysis between ICAT variants and human judgments (see Table 2) reveals strong performance across most evaluation methods. Using Llama-3.1-70B as the coverage model, ICAT-S achieves the strongest correlations (Pearson’s $\rho$ = 0.422, p < 0.01; Spearman’s $\rho$ = 0.446, p < 0.01). Interestingly, ICAT-A demonstrates better performance than ICAT-M despite not requiring ground-truth aspects. While ICAT-A seems to have much weaker correlation than ICAT-S, it is worth noting that in practice, we observed the automatically generated topics to be higher quality and more exhaustive. Using a more comprehensive set of ground-truth topics, ICAT-A would probably demonstrate higher correlation with human judgements.
Comparison with Traditional Metrics for Text Generation.
To demonstrate the advantages of ICAT over traditional text generation evaluation metrics for assessing topic coverage, we compared our approach with widely-used metrics including BLEU, ROUGE, METEOR, and BERTScore. For this comparison, we concatenated the text of all ground-truth topics as a reference and evaluated the LLM responses against this reference. Table 4 presents the correlation results with human-annotated coverage scores.
Among traditional metrics, only BERTScore-recall shows statistically significant correlation with ground-truth coverage scores (Pearson’s $\rho$ = 0.291, p < 0.001). However, this correlation is still much lower than ICAT-S. Notably, n-gram based metrics (BLEU, ROUGE variants) show negative or near-zero correlations, demonstrating their inadequacy for evaluating topic coverage in long-form text. These results confirm that traditional metrics, designed primarily for surface-level text similarity, fail to capture the semantic coverage of diverse aspects that ICAT effectively measures.
Metric Pearson $\rho$ p-value Spearman $\rho$ p-value Kendall $\tau$ p-value
BERTScore-precision -0.003 0.966 0.048 0.499 0.019 0.709
BERTScore-recall 0.291 0.000 0.313 0.000 0.226 0.000
BERTScore-F1 0.120 0.090 0.188 0.008 0.126 0.013
METEOR 0.027 0.704 0.100 0.160 0.071 0.164
ROUGE-1 -0.058 0.415 0.016 0.819 0.014 0.785
ROUGE-2 -0.090 0.204 -0.032 0.655 -0.018 0.725
ROUGE-L -0.090 0.206 0.020 0.781 0.012 0.808
ROUGE-Lsum -0.090 0.206 0.020 0.781 0.012 0.808
BLEU -0.099 0.165 -0.059 0.409 -0.046 0.413
Table 4: Correlation of traditional text generation metrics with human-annotated topic coverage scores. Only BERTScore-recall shows statistically significant correlation, though still lower than ICAT variants.
@@
Our experimental results reveal distinct patterns in how different LLMs balance factuality and coverage (see Table 3). When using corpus-based retrieval, Llama-3-70B demonstrates superior Coverage Score (0.451), while GPT-4 and Mixtral-8x22B show comparable factuality scores (0.343 and 0.344, respectively). However, Mixtral exhibits notably lower Coverage Score (0.370) compared to GPT-4’s 0.416, resulting in lower overall $\text{ICAT}_{1}$ scores (0.297 vs 0.327). Notably, when employing web-based retrieval, we observe substantially higher factuality scores across all models. This significant improvement can be attributed to the broader knowledge base available through web search, allowing more claims to be successfully grounded. Coverage scores also show improvement with web-based retrieval, though the increase is more modest. This suggests that while web search enables better fact verification, the comprehensiveness of aspect coverage is more dependent on the model’s capabilities than the retrieval source.
@@
Component-level Evaluation.
To understand the contribution of each component in our framework, we conducted detailed evaluations of the atomic claim generation, topic generation, and coverage assessment modules. A comprehensive analysis of each component’s performance is provided in Appendix A.5.